# Leveraging LLMs for Top-Down Sector Allocation in Automated Trading

### v1

@@ Abstract
This paper introduces a methodology leveraging Large Language Models (LLMs) for sector-level portfolio allocation through systematic analysis of macroeconomic conditions and market sentiment. Our framework emphasizes top-down sector allocation by processing multiple data streams simultaneously, including policy documents, economic indicators, and sentiment patterns. Empirical results demonstrate superior risk-adjusted returns compared to traditional cross momentum strategies, achieving a Sharpe ratio of 2.51 and portfolio return of 8.79% versus -0.61 and -1.39% respectively. These results suggest that LLM-based systematic macro analysis presents a viable approach for enhancing automated portfolio allocation decisions at the sector level.
@@ 1 Introduction
Macroeconomic factors fundamentally shape investment decision-making, influencing stock valuations through both economic channels and market sentiment (Jareño and Negrut, 2016; Jabeen et al., 2022; Ma et al., 2024). A top-down investment approach, based on macroeconomic analysis, enables investors to anticipate broad market trends, optimize portfolio allocation across economic cycles, and identify systematic risks before they manifest in individual equity prices.
These macroeconomic forces generate heterogeneous sectoral responses; for instance, inflationary pressures typically constrain manufacturing growth while benefiting services and agricultural sectors (Chaudhry et al., 2013). Moreover, hawkish Federal Reserve communications not only elevate borrowing costs and alter corporate decision-making but also intensify investor concerns about financially vulnerable firms, amplifying negative sentiment (Cieslak and McMahon, 2023). This sentiment-driven reaction often precipitates market movements that exceed what fundamentals alone would otherwise suggest. Key macroeconomic variables—including monetary policy, inflation dynamics, and economic growth trajectories—propagate through the financial system, influencing both firms’ cost of capital and aggregate consumption patterns.
@@
Given that the task of constructing a portfolio based on sector allocation is one that requires piecing together information from different sources, this paper seeks to address the following Research Question – How can LLMs be leveraged to enhance top-down sector allocation strategies by integrating macroeconomic analysis with market sentiment for automated portfolio construction?
@@
To respond to this question, we introduce a novel LLM-based methodology that emphasizes top-down sector allocation through integrated analysis of macroeconomic and market sentiment data. While existing approaches incorporate various data sources including sentiment and fundamentals, they typically focus on bottom up analysis of individual securities. Our framework leverages LLMs to systematically process and synthesize multiple data streams simultaneously (including policy documents, economic indicators, and sentiment patterns), enabling dynamic adjustment of sector allocations based on market conditions. By automating the extraction and interpretation of these macro financial relationships, our framework enhances the responsiveness of sector allocation strategies through a predominantly top-down lens. This provides a more systematic approach to capturing sentiment driven price movements, offering new insights into sector level portfolio construction that complement traditional security level analysis.
@@
Results obtained through backtesting indicate that our sector allocation strategy significantly outperforms traditional cross momentum approaches. Over the testing period, our methodology achieved a positive return of 8.79% and a Sharpe ratio of 2.51, compared to a loss of 1.39% and a Sharpe ratio of -0.61 for the cross momentum strategy. These results suggest that our top-down LLM-based approach effectively captures sector-level opportunities while maintaining strong risk-adjusted performance, demonstrating the potential value of incorporating systematic macro and sentiment analysis in portfolio allocation decisions.
@@ 2.3.2 LLM as a Trader - Reflection Driven
Similarly, FinAgent (Zhang et al., 2024) extends this approach with a multimodal module that processes numeric, text, and image data. It enhances decision-making by incorporating technical indicators like MACD and RSI, along with analyst guidance, demonstrating superior backtesting performance over FinMem.
@@ 2.4 Finance specific Aspect based sentiment analysis
Traditional sentiment analysis in finance has historically focused on broad positive/negative classifications of text towards stocks, but markets and financial decisions require more nuanced understanding. Aspect based Sentiment Analysis (ABSA) has emerged as a crucial advancement by enabling the extraction and analysis of sentiment toward specific aspects or features within financial text Du et al. (2024). Previous works of finance specific ABSA  (Ong et al., 2023) highlight the separation of sentiment across different financial aspects for explainable sentiment, allowing for assessment of sentiment towards specific risk factors or market aspects.
@@ 7 Limitations and Future Directions
Several promising directions exist for future research. First, expanding the model’s data inputs to include company fundamental data, particularly quarterly earnings reports, could reveal whether incorporating bottom up financial metrics alongside our top-down approach yields meaningful improvements in portfolio performance (Ong et al., 2025).
Second, the ranking mechanism could be enhanced through the implementation of LTR algorithms as mentioned in Section 2.1 or Reinforcement Learning techniques. These approaches have demonstrated success in improving cross sectional strategies, suggesting potential benefits for sector allocation. By incorporating historical ranking performance as feedback, these methods could dynamically optimize allocation decisions and enhance the model’s adaptability to changing market conditions.
Additionally, integrating FINMEM’s (Yu et al., 2024) architecture could enhance the model’s ability to process multi-timeframe financial data while maintaining important information over extended periods, potentially capturing longer term patterns and relationships that human traders might overlook.
Lastly, financial Explainable Artificial Intelligence techniques (Yeo et al., 2023) such as counterfactual explanations, can be used to evaluate performance in alternative scenarios. These methods provide insight into LLM decision-making and support regulatory compliance.
@@ 8 Conclusion
In this paper, we introduced a novel methodology leveraging Large Language Models for top-down sector allocation through integrated analysis of macroeconomic and market sentiment data. Our approach demonstrates that LLMs can effectively process and synthesize multiple data streams simultaneously, enabling dynamic sector allocation adjustments based on market conditions. The empirical results validate the effectiveness of this approach, with our strategy achieving an 8.79% return and a Sharpe ratio of 2.51, significantly outperforming traditional cross momentum approaches which recorded a -1.39% return and -0.61 Sharpe ratio during the same period.
These results suggest that systematic top-down analysis using LLMs offers a viable approach for enhancing portfolio allocation decisions. The framework’s ability to process unstructured data and identify macro financial relationships represents a step forward in automating complex market analysis tasks. As the field of quantitative investing continues to evolve, our methodology demonstrates the potential for LLMs to complement existing investment strategies by providing a more systematic approach to sector level portfolio construction.
@@ Acknowledgments
Gianmarco Mengaldo gratefully acknowledges the support of the MOE Tier 1 Startup project, under grant number #22-3565-A0001-1.

### v3

@@ Abstract
This paper introduces a methodology leveraging Large Language Models (LLMs) for sector-level portfolio allocation through systematic analysis of macroeconomic conditions and market sentiment. Our framework emphasizes top-down sector allocation by processing multiple data streams simultaneously, including policy documents, economic indicators, and sentiment patterns. Empirical results demonstrate superior risk-adjusted returns compared to traditional cross-momentum strategies, achieving a Sharpe ratio of 2.51 and portfolio return of 8.79% versus -0.61 and -1.39% respectively. These results suggest that LLM-based systematic macro analysis presents a viable approach for enhancing automated portfolio allocation decisions at the sector level.
@@ 1 Introduction
Macroeconomic factors fundamentally shape investment decision-making, influencing stock valuations through both economic channels and market sentiment (Jareño and Negrut, 2016; Jabeen et al., 2022; Ma et al., 2024). A top-down investment approach, based on macroeconomic analysis, enables investors to anticipate broad market trends, optimize portfolio allocation across economic cycles, and identify systematic risks before they manifest in individual equity prices. These macroeconomic forces generate heterogeneous sectoral responses; for instance, inflationary pressures typically constrain manufacturing growth while benefiting services and agricultural sectors (Chaudhry et al., 2013). Moreover, hawkish Federal Reserve communications not only elevate borrowing costs and alter corporate decision-making but also intensify investor concerns about financially vulnerable firms, amplifying negative sentiment (Cieslak and McMahon, 2023).
@@
This sentiment-driven reaction often precipitates market movements that exceed what fundamentals alone would otherwise suggest. Key macroeconomic variables—including monetary policy, inflation dynamics, and economic growth trajectories—propagate through the financial system, influencing both firms’ cost of capital and aggregate consumption patterns. Given that the task of constructing a portfolio based on sector allocation is one that requires piecing together information from different sources, this paper seeks to address the following Research Question – How can LLMs be leveraged to enhance top-down sector allocation strategies by integrating macroeconomic analysis with market sentiment for automated portfolio construction?
@@
To respond to this question, we introduce a novel LLM-based methodology that emphasizes top-down sector allocation through integrated analysis of macroeconomic and market sentiment data. While existing approaches incorporate various data sources including sentiment and fundamentals, they typically focus on bottom-up analysis of individual securities. Our framework leverages LLMs to systematically process and synthesize multiple data streams simultaneously (including policy documents, economic indicators, and sentiment patterns), enabling dynamic adjustment of sector allocations based on market conditions. By automating the extraction and interpretation of these macro financial relationships, our framework enhances the responsiveness of sector allocation strategies through a predominantly top-down lens. This provides a more systematic approach to capturing sentiment-driven price movements, offering new insights into sector-level portfolio construction that complement traditional security-level analysis.
@@
Results obtained through backtesting indicate that our sector allocation strategy significantly outperforms traditional cross-momentum approaches. Over the testing period, our methodology achieved a positive return of 8.79% and a Sharpe ratio of 2.51, compared to a loss of 1.39% and a Sharpe ratio of -0.61 for the cross-momentum strategy. These results suggest that our top-down LLM-based approach effectively captures sector-level opportunities while maintaining strong risk-adjusted performance, demonstrating the potential value of incorporating systematic macro and sentiment analysis in portfolio allocation decisions.
@@ 2.3.2 LLM as a Trader - Reflection Driven
Similarly, FinAgent (Zhang et al., 2024) extends this approach with a multimodal module that processes numeric, text, and image data. It enhances decision making by incorporating technical indicators like MACD and RSI, along with analyst guidance, demonstrating superior backtesting performance over FinMem.
@@ 2.4 Finance specific Aspect based sentiment analysis
Traditional sentiment analysis in finance has historically focused on broad positive/negative classifications of text towards stocks, but markets and financial decisions require more nuanced understanding. Aspect based Sentiment Analysis (ABSA) has emerged as a crucial advancement by enabling the extraction and analysis of sentiment toward specific aspects or features within financial text Du et al. (2024); Malandri et al. (2018); Ma et al. (2023). Previous works of finance-specific ABSA (Ong et al., 2023) highlight the separation of sentiment across different financial aspects for explainable sentiment, allowing for assessment of sentiment towards specific risk factors or market aspects.
@@ 7 Limitations and Future Directions
Several promising directions exist for future research. First, expanding the model’s data inputs to include company fundamental data, particularly quarterly earnings reports, could reveal whether incorporating bottom-up financial metrics alongside our top-down approach yields meaningful improvements in portfolio performance (Ong et al., 2025).
Second, the ranking mechanism could be enhanced through the implementation of LTR algorithms as mentioned in Section 2.1 or Reinforcement Learning techniques. These approaches have demonstrated success in improving cross-sectional strategies, suggesting potential benefits for sector allocation. By incorporating historical ranking performance as feedback, these methods could dynamically optimize allocation decisions and enhance the model’s adaptability to changing market conditions.
Additionally, integrating FINMEM’s (Yu et al., 2024) architecture could enhance the model’s ability to process multi-timeframe financial data while maintaining important information over extended periods, potentially capturing longer-term patterns and relationships that human traders might overlook.
Lastly, financial Explainable Artificial Intelligence techniques (Yeo et al., 2025) such as counterfactual explanations, can be used to evaluate performance in alternative scenarios. These methods provide insight into LLM decision making and support regulatory compliance.
@@ 8 Conclusion
In this paper, we introduced a novel methodology leveraging Large Language Models for top-down sector allocation through integrated analysis of macroeconomic and market sentiment data. Our approach demonstrates that LLMs can effectively process and synthesize multiple data streams simultaneously, enabling dynamic sector allocation adjustments based on market conditions. The empirical results validate the effectiveness of this approach, with our strategy achieving an 8.79% return and a Sharpe ratio of 2.51, significantly outperforming traditional cross-momentum approaches which recorded a -1.39% return and -0.61 Sharpe ratio during the same period.
These results suggest that systematic top-down analysis using LLMs offers a viable approach for enhancing portfolio allocation decisions. The framework’s ability to process unstructured data and identify macro financial relationships represents a step forward in automating complex market analysis tasks. As the field of quantitative investing continues to evolve, our methodology demonstrates the potential for LLMs to complement existing investment strategies by providing a more systematic approach to sector-level portfolio construction.
@@ Acknowledgments
Gianmarco Mengaldo gratefully acknowledges the support of the MOE Tier 1 Startup project, under grant number #22-3565-A0001-1. The research is also supported by MOE Academic Research Fund Tier 2 (STEM RIE2025 Award MOE-T2EP20123-0005) and RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore.